5. Lo spun{} equal to all of F? Til
Answer #1

10:1:11:13 A = [ 1 1 0 if Talto Then Otherwise not it is span of R3 LAI 3 1 1 1 0 1 (1-2)-1 (0-1) = -1+1=0 not spanned of R